#6.) Given f(t)=-2:+8, Ost<4, f(t+4)= f(t). Find F(s)=L{f(t)} of the Periodic Function.
Answer #1

Please find attached solution..

Given that f(t)=-1++8, ostia and f(t+4)=f(t) :f(t) to a periodic function with the period T=4. We know that, = f(t) estat (5)Date T -45 & é डप + -88-1 52 52 T= 안 -45-4 २ -पड 4 e S + 1 1- 8 (1) 52 fromo) L { (j - T -45 -पड -45 -1 S + 4e -65 ] د کام 8